# Stable Offline Value Function Learning with Bisimulation-based Representations

### v2

@@ 4.2 Analyzing Fundamental Properties of the Learned Representations
(b) Realizability Error
(c) Orthogonality
Figure 2: Evaluation of basic representation properties on Garnet mdps with $40$ state-actions vs. output dimension $d$. Figure 2(a): Fractions of runs out of $30$ trials that resulted in spectral radius of $(\mathbb{E}[\Phi^{\top}\Phi])^{-1}(\gamma\mathbb{E}[\Phi^{\top}P^{{\pi_{e}}}\Phi])$ to be less than $1$; higher is better. Figure 2(b): Realizability error; lower is better. Figure 2(c): Pearson correlation between orthogonality between pairs of latent features vs. their corresponding absolute $q^{{\pi_{e}}}$ action-value difference; higher is better. All results are averaged over $30$ trials and the shaded region is the $95\%$ confidence interval.
@@
Realizability.
A basic criterion for learning $q^{\pi_{e}}$ is realizability. That is, we want ${\epsilon}:=\|\Phi\hat{w}-q^{{\pi_{e}}}\|^{2}_{2}$, where $\hat{w}:=\operatorname*{arg\,min}_{w}\|\Phi w-q^{{\pi_{e}}}\|_{2}^{2}$, to be low. In our experiments, we compute ${\epsilon}$ and plot it as a function of $d$ in Figure 2(b). A critical message from our results is that stability and realizability do not always go hand-in-hand. While bcrl-exp-na has favorable spectral radius properties (Figure 2(a)), it has poor realizability, which will negatively affect its ope accuracy. krope, on the other hand, has favorable stability and realizability properties up till $d=30$. When $d\geq 40$, the realizability error is 00 for all algorithms since the subspace spanned by $\Phi$ is large enough to contain the true action-value function (Ghosh & Bellemare, 2020). While the realizability error is 00 for $d\geq 40$, the representations can be unstable (Figure 2(a)).
Generalization.
Finally, we say that the representations have generalized well when state-actions that have similar $q^{\pi_{e}}$ values are close to each other in the representation space (Lyle et al., 2022). We assess generalization by measuring the orthogonality: $1-\frac{|\langle\phi(s_{1},a_{1}),\phi(s_{2},a_{2})\rangle|}{\|\phi(s_{1},a_{1})\|\|\phi(s_{2},a_{2})\|}$ (Wang et al., 2024) between every state-action pair, $(s_{1},a_{1};s_{2},a_{2})$, and the absolute action-value difference: $|q^{\pi_{e}}(s_{1},a_{1})-q^{\pi_{e}}(s_{2},a_{2})|$. We then compute the Pearson correlation between these values for every pair and plot the correlation for each $d$ in Figure 2(c). A correlation coefficient close to $1$ indicates that the representations generalize well. We find that krope representations satisfy this property almost perfectly since it specifically tries to learn representations such that state-action pairs with similar values under ${\pi_{e}}$ are similar. We observe that the other algorithms typically have zero or even negative correlation. A negative correlation indicates that state-actions with different action-values may be similar in latent space, which can result in higher realizability error (Figure 2(b)). A near-zero correlation but low realizability error such as in the case of fqe implies that accurate offline value prediction is still possible but that it generalizes poorly, which may slow down convergence to its ope solution (Lyle et al., 2022).
@@ 4.3 Offline Policy Evaluation
In this set of experiments, we conduct experiments on four DM control environments to answer our second empirical question: whether krope representations lead to stable and low msve? We also evaluate bcrl-na, which is bcrl without the exploration maximization regularizer. To stabilize training for all algorithms, we use wide neural networks with layernorm (Gallici et al., 2024; Ota et al., 2021). Note that while wide networks and layernorm stabilize training, they may not lead to stable lspe under the linear evaluation protocol. During representation learning, we periodically evaluate the learned representations for ope using lspe. The corresponding (normalized) squared value errors are presented in Figure 3.
@@
(b) CheetahRun
(c) FingerEasy
(d) WalkerStand
Figure 3: Normalized squared value error achieved by lspe when using a particular representation vs. representation training epochs. lspe estimates are computed every $10$ epochs. Results are averaged over $20$ trials and the shaded region is the $95\%$ confidence interval. Lower and less erratic is better.
In general, we find that krope representations lead to low and stable msve. On the other hand, we find that the other auxiliary tasks inconsistently produce stable ope estimates across all environments. For example, the performance of the abs-dr3 and beer regularizer suggests that explicitly trying to increase the rank of the features of the penultimate layer may hurt stability, and even if the ope error is stable, it can hurt accuracy. We also make a similar observation for bcrl. However, in this case, we attribute poor performance to difficulty in optimizing the bcrl objective. In fact, in Figure 4(c), we will see that bcrl is sensitive to hyperparameter tuning. We also observe results consistent with a known result that bcrl-exp-na performs better than bcrl-na indicating the known result that exploration maximization of the covariance matrix helps produce stable representations (Chang et al., 2022). Finally, while fqe achieves lower ope error than krope on WalkerStand, it is very unstable on CartPoleSwingUp and Finger Easy, which motivates the need to shape the representations for stable and accurate ope. We note that in the WalkerStand instance, the higher error of krope is unsurprising since Lemma 3 suggests that krope may lose realizability (see Appendix B.1). We refer the reader to Appendix C.1 for details of each algorithm. We also conduct the same experiment on $9$ d4rl datasets and reach the similar conclusions (see Appendix C.3.2).
(a) Feature Co-adaptation
(b) Condition Number
(c) Hyperparameter Sensitivity
Figure 4: Stability-related metrics to understand the properties of krope on CartPoleSwingUp. Values are plotted as a function of training epochs. All results are averaged over $20$ trials and shaded region is the $95\%$ confidence interval. For hyperparameter sensitivity, larger area under the curve is better.
4.3.1 Analyzing Stability-Related Metrics
While krope performs well on the downstream task of ope, it is important to analyze upstream stability-related metrics during the course of learning. These results give insight into the properties of the learned representations. We present the results in Figure 4 for CartPoleSwingUp and defer the remaining results to the Appendix C.
Feature co-adaptation. The feature-co-adaptation metric, i.e., $\sum_{(s,a,s^{\prime})\in\mathcal{D},a^{\prime}\sim{\pi_{e}}}\phi(s,a)^{\top}\phi(s^{\prime},a^{\prime})$ was shown to correlate with instability as high values can indicate representation collapse (Kumar et al., 2021). From Figure 4(a), we find that krope mitigates representation collapse of fqe (lowers feature co-adaptation from $10^{6}$ to $10^{5}$). Its relatively moderate value indicates that it is able to balance generalization and maintain distinctness between state-action representations. While low co-adaptation is preferred, extremely low values such as those achieved by abs-dr3 do not necessarily lead to accurate ope (see Figure 3). These results tell us which algorithms are more prone to representation collapse.
Condition number. Another metric that we analyze is the condition number of the covariance matrix $M:=\mathbb{E}[\Phi^{\top}\Phi]$. A low condition number indicates numerical stability of the learning algorithm, i.e., solutions recovered by lspe (see Equation (1)) or general td algorithms (Asadi et al., 2024) are less likely to change with small perturbations to $M$ (Chang et al., 2022). From Figure 4(b), we find that krope achieves lower condition number than fqe (from $10^{4}$ to $10^{1}$). As excepted, we observe that bcrl-exp-na achieves an even lower condition since it explicitly optimizes for a well-conditioned covariance matrix ($10^{4}$ to $10^{3}$). While there is no ideal value, these results indicate the reliability of the representations in enabling td-based algorithms to recover their value function solution. We observe qualitatively similar results for other environments (see Appendix C).
Hyperparameter Sensitivity. In ope, hyperparameter tuning can be challenging since it may be infeasible to get access to ground truth performance of ${\pi_{e}}$ (Fu et al., 2021). Therefore, we prefer algorithms that are robust to hyperparameter tuning, i.e, they reliably produce accurate ope estimates for a wide range of hyperparameters. In Figure 4(c), we present the performance profile for each algorithm across all hyperparameter combinations and all trials (Agarwal et al., 2021). We tune the hyperparameters discussed in Appendix C.1. We find that $100\%$ krope runs across all instances produce msve $\leq 1$, which is not the case with other algorithms.
@@ 5 Limitations and Future Work
In this section, we discuss limitations and future work. A shortcoming of our work is that krope’s learning algorithm is susceptible to instability since it is a semi-gradient method (Sutton & Barto, 2018). Moreover, its fixed-point optimization means it does not solve any objective function (Feng et al., 2019). In our work, we employed commonly-used techniques such as layernorm and wide neural networks to mitigate instability (Ota et al., 2021; Gallici et al., 2024). While these techniques potentially side-step the issue, the consequences of a semi-gradient method may still exist. In Appendix C.3.1, we present an empirical analysis to gauge when krope’s learning algorithm may be unstable. We find that while individual off-policy transitions can determine the instability of fixed-point and semi-gradient algorithms such as fqe, pairs of off-policy transitions can determine krope’s instability. Since we are unlikely to have control over the distribution over pairs of transitions in practice, we need to resort to fundamental changes to the algorithm. One potential change is based on that by Feng et al. (2019). Their key insight is to leverage the Legendre-Fenchel transformation from optimization theory and replace the fixed-point loss function of semi-gradient methods with an equivalent expression that avoids semi-gradient learning (Rockafellar & Wets, 1998). However, a drawback with this approach is that the new learning objective is a minimax procedure, which can be challenging to optimize in practice. In future work, we will explore the viability of this approach to design a provably convergent version of krope.
@@ 6 Conclusion
Acknowledgements
The authors thank Abhinav Narayan Harish, Adam Labiosa, Andrew Wang, Anshuman Senapati, David Abel, Kyle Domico, Ishan Durugkar, and Prabhat Nagarajan for feedback on earlier versions of this draft, and thank Fanghui Liu for inspiring discussions on kernel methods. Y. Chen is partially supported by NSF CCF-1704828 and NSF CCF-2233152.
@@ A.1 Bisimulation Metrics
In this section, we present background information on bisimulations and its associated metrics. Our proposed representation learning algorithm is a bisimulation-based algorithm. Bisimulation abstractions are those under which two states with identical reward functions and that lead to identical groups of next states under any action are classified as similar (Ferns et al., 2004; 2011; Ferns & Precup, 2014). Bisimulations are the strictest forms of abstractions. In practice, the exact bisimulation criterion is difficult to satisfy computationally and statistically. A more relaxed version of this notion is the $\pi$-bisimulation metrics. These metrics capture the similarity between two states such that two states are considered similar if they have identical expected reward functions and expected transitions to identical groups of next states under a fixed policy $\pi$ (Castro, 2019).
@@
Using the above metrics, prior works have introduced several representation learning algorithms to learn representations such that the distance between representations in latent space model the above distance metrics (Castro et al., 2022; 2023; Zhang et al., 2021; Kemertas & Aumentado-Armstrong, 2021; Pavse & Hanna, 2023a).
@@ A.2 Reproducing Kernel Hilbert Spaces
The core usage of the rkhs is to precisely characterize the nature of the krope kernel. In practice, we deal with neural network representations, which are embedded in Euclidean space. Therefore, our goal is to learn representations in Euclidean space that approximate the properties of representations in the rkhs. For more details on the rkhs, we refer readers to Castro et al. (2023) and Gretton et al. (2012).
@@ Appendix B Theoretical Results
In this section, we present the proofs of our main and supporting theoretical results. The first set of proofs in Section B.1 show that krope is a valid operator. While new to our work, the proofs follow those by Castro et al. (2023). The next set of proofs in Section B.2 prove the stability of krope representations and are novel to our work. For presentation purposes, it will often be convenient to refer to a state-action pair as $x\in\mathcal{X}$ instead of $(s,a)$.
@@
We now present the proofs demonstrating the validity of the krope operator. All the proofs in this sub-section model those by Castro et al. (2023). The primary difference is that our operator is for state-actions instead of states.
@@
Before presenting the remaining results, we define the distance metric $d_{\text{{krope}}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ induced by the krope kernel $k^{{\pi_{e}}}$ as follows:
@@ B.2.1 Supporting Theoretical Results
We first define ${\pi_{e}}$-bisimulation Castro (2019). Note that Castro (2019) considered only state abstractions, while we consider state-action abstractions. $\phi$ is considered a ${\pi_{e}}$-bisimulation abstraction if it induces a mapping between $\mathcal{X}$ and $\mathcal{X}^{\phi}$ such that for any $x,y\in\mathcal{X}$ such that $x,y\in\phi(x)$, we have:
@@
Once we have $\phi$, we consider the following to show Bellman completeness. Our proof closely follows the proof of Proposition 20 from Chen & Jiang (2019). First recall the definition of Bellman completeness from Definition 8: $\forall f\in\mathcal{F},\forall\mathcal{T}^{{\pi_{e}}}f\in\mathcal{G}$, $\sup_{f\in\mathcal{F}}\inf_{g\in\mathcal{G}}\|g-\mathcal{T}^{{\pi_{e}}}f\|_{\infty}=0$. Given that the smallest value $\forall f\in\mathcal{F},\forall\mathcal{T}^{{\pi_{e}}}f\in\mathcal{G}$, $\sup_{f\in\mathcal{F}}\inf_{g\in\mathcal{G}}\|g-\mathcal{T}^{{\pi_{e}}}f\|_{\infty}$ can take on is zero, we will prove our claim by showing that $\forall f\in\mathcal{F},\forall\mathcal{T}^{{\pi_{e}}}f\in\mathcal{G}$, $\sup_{f\in\mathcal{F}}\inf_{g\in\mathcal{G}}\|g-\mathcal{T}^{{\pi_{e}}}f\|_{\infty}$ is upper-bounded by zero when $\phi$ is a ${\pi_{e}}$-bisimulation.
@@
Proposition 2.21 from Paulsen & Raghupathi (2016) states that any kernel $k$ is positive semidefinite if it takes the form: $k(a,b)=\min\{a,b\}$ where $a,b\in[0,\infty)$.
@@
From Gretton et al. (2012), we have the definition of MMD between two probability distributions $p,q$ given kernel $k$:
@@
By Proposition 2.21 in Paulsen & Raghupathi (2016), the matrix $K$ is positive semidefinite, so we only need to show that $K$ is full rank. WLOG assume that $0<x_{1}<x_{2}<\cdots<x_{n}$. We prove by induction on $n$. The base case with $n=1$ clearly holds. Suppose the claim holds for $n-1$ numbers. Now consider $n$ numbers. Let $\alpha=(\alpha_{1},\ldots,\alpha_{n})^{\top}\in^{n}$. It suffices to show that $K\alpha=0$ implies $\alpha=0$. We write $K$ in block matrix form as

### v3

@@ 4.2 Analyzing Stability and $q^{\pi_{e}}$-Consistency Properties
(b) Orthogonality
Figure 2: Evaluation of basic representation properties on Garnet mdps with $40$ state-actions vs. output dimension $d$. Figure 2(a): Fractions of runs out of $30$ trials that resulted in spectral radius of $(\mathbb{E}[\Phi^{\top}\Phi])^{-1}(\gamma\mathbb{E}[\Phi^{\top}P^{{\pi_{e}}}\Phi])$ to be less than $1$; higher is better. Figure 2(b): Pearson correlation between orthogonality between pairs of latent features vs. their corresponding absolute $q^{{\pi_{e}}}$ action-value difference; higher is better. All results are averaged over $30$ trials and the shaded region is the $95\%$ confidence interval.
@@
$q^{\pi_{e}}$-Consistency.
We say that the representations have maintained $q^{\pi_{e}}$-consistency well when state-actions that have similar $q^{\pi_{e}}$ values are close to each other in the representation space (Lan et al., 2021). We assess this property by first measuring the orthogonality: $1-\frac{|\langle\phi(s_{1},a_{1}),\phi(s_{2},a_{2})\rangle|}{\|\phi(s_{1},a_{1})\|\|\phi(s_{2},a_{2})\|}$ (Wang et al., 2024) between every state-action pair, $(s_{1},a_{1};s_{2},a_{2})$, and the absolute action-value difference: $|q^{\pi_{e}}(s_{1},a_{1})-q^{\pi_{e}}(s_{2},a_{2})|$. We then compute the Pearson correlation between these values for every pair and plot the correlation for each $d$ in Figure 2(b). A correlation coefficient close to $1$ indicates that the representations maintain $q^{\pi_{e}}$-consistency well. We find that krope representations satisfy this property almost perfectly since it specifically tries to learn representations such that state-action pairs with similar values under ${\pi_{e}}$ are similar. We observe that the other non-bisimulation-based algorithms typically have zero or even negative correlation. A negative correlation indicates that state-actions with different action-values may have similar representations in latent space, which can complicate value function learning.
@@ 4.3 Offline Policy Evaluation
In this set of experiments, we evaluate the algorithms for ope on $13$ datasets: $4$ dm control and $9$ d4rl datasets. We also evaluate bcrl-na, which is bcrl without the exploration maximization regularizer. To stabilize training for all algorithms, we use wide neural networks with layernorm (Gallici et al., 2024; Ota et al., 2021). Note that while wide networks and layernorm stabilize training, they may not lead to stable lspe under the linear evaluation protocol. We report the final (normalized) squared value errors when the learned representations are used with lspe for ope. The results are presented in Table 1 (see Appendix C.3.2 for full learning curves).
Algorithm
Dataset (dmc) fqe bcrl+exp bcrl beer dr3 dbc rope krope (ours)
CartPoleSwingUp Div. $2.0\pm 1.6$ $2.2\pm 0.8$ Div. $0.9\pm 0.0$ Div. $0.2\pm 0.1$ $0.0\pm 0$ 0.0
CheetahRun $0.0\pm 0$ 0.0 $0.3\pm 0.2$ $0.8\pm 0.3$ $0.0\pm 0$ 0.0 $0.4\pm 0.0$ Div. Div. $0.0\pm 0$ 0.0
FingerEasy Div. $0.6\pm 0.1$ $0.8\pm 0.2$ Div. $0.9\pm 0.0$ Div. $0.1\pm 0.0$ $0.6\pm 0.0$
WalkerStand $0.0\pm 0$ 0.0 $0.2\pm 0.2$ $0.2\pm 0.1$ $1.9\pm 3.6$ $0.1\pm 0.0$ Div. $0.2\pm 0.0$ $0.0\pm 0$ 0.0
Dataset (d4rl) fqe bcrl+exp bcrl beer dr3 dbc rope krope (ours)
cheetah random $0.9\pm 0.0$ Div. Div. $0.9\pm 0.0$ $0.9\pm 0.0$ $0.9\pm 0.0$ $1.0\pm 0.0$ $1.0\pm 0.0$
cheetah medium Div. Div. $0.2\pm 0.2$ Div. Div. Div. $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
cheetah med-expert Div. $0.2\pm 0.1$ $0.3\pm 0.1$ Div. Div. Div. $0.1\pm 0.0$ $0.0\pm 0$ 0.0
hopper random Div. Div. Div. Div. $0.8\pm 0.0$ Div. Div. $0.1\pm 0.0$
hopper medium Div. Div. Div. Div. Div. Div. Div. Div.
hopper med-expert Div. Div. Div. Div. $0.6\pm 0.0$ Div. $0.0\pm 0$ 0.0 $0.0\pm 0$ 0.0
walker random Div. Div. Div. Div. $1.0\pm 0.0$ Div. Div. $0.5\pm 0.1$
walker medium Div. Div. Div. Div. Div. Div. Div. Div.
walker med-expert Div. $1.3\pm 0.4$ $2.6\pm 2.1$ Div. $6.6\pm 11.6$ Div. $0.1\pm 0.0$ Div.
Table 1: Final normalized squared value error achieved by lspe with different representation learning algorithms. Results are averaged over $20$ trials and the variation is the $95\%$ confidence interval. Lower and less erratic is better. Values are rounded to single place decimal. When an algorithm’s error is $\geq 10$, we label it as diverged (Div.). Bolded and highlighted error indicates lowest error among baselines.
In general, we find that krope representations lead to low and stable msve. On the other hand, we find that the other auxiliary tasks inconsistently produce stable ope estimates across all environments. Compared to the other bisimulation-based algorithms (dbc and rope), krope consistently achieves lower error. In all cases, dbc seems to be unreliable for ope. rope, which was the previous state-of-the-art bisimulation-based algorithm for ope, also achieves low error. While rope is competitive with krope, krope generally outperforms rope and has one less hyperparameter. The performance of the abs-dr3 and beer regularizer suggests that explicitly trying to increase the rank of the features of the penultimate layer may hurt stability, and even if the ope error is stable, it can hurt accuracy. We also make a similar observation for bcrl. However, in this case, we attribute poor performance to difficulty in optimizing the bcrl objective. In fact, in Figure 3(a), we will see that bcrl is sensitive to hyperparameter tuning. We also observe results consistent with a known result that bcrl-exp-na achieves lower error than bcrl-na indicating the known result that exploration maximization of the covariance matrix helps produce stable representations (Chang et al., 2022). While fqe achieves low error on WalkerStand and CheetahRun, it is very unstable on the other datasets, which motivates the need to shape the representations for stable and accurate ope. Note that in practice, krope may still diverge, as it did it in $3$ cases, because the learning objective in Equation (4) is still a semi-gradient learning algorithm (see discussion in Section 5 and Appendix C.3.1).
@@
(b) WalkerStand
Figure 3: Hyperparameter sensitivity plots on CartPoleSwingUp and WalkerStand. Results are over $20$ trials for each hyperparameter configuration and shaded region is the $95\%$ confidence interval. Larger area under the curve is better.
Hyperparameter Sensitivity.
In ope, hyperparameter tuning can be challenging since it may be infeasible to get access to ground truth performance of ${\pi_{e}}$ (Fu et al., 2021). Therefore, we prefer algorithms that are robust to hyperparameter tuning, i.e, they reliably produce accurate ope estimates for a wide range of hyperparameters. In Figure 3, we present the performance profile for each algorithm across all hyperparameter combinations and all trials (Agarwal et al., 2021). We tune the hyperparameters discussed in Appendix C.1. We find that $100\%$ krope runs across all instances produce msve $\leq 1$, which is not the case with other algorithms. This result indicates the reliability of using krope for ope over other algorithms including other bisimulation-based algorithms.
@@ 5 Limitations and Future Work
In this section, we discuss limitations and future work. A shortcoming of our work is that krope’s learning algorithm is susceptible to instability since it is a semi-gradient method (Sutton & Barto, 2018). Moreover, its fixed-point optimization means it does not solve any objective function (Feng et al., 2019). In our work, we employed commonly-used techniques such as layernorm and wide neural networks to mitigate instability (Ota et al., 2021; Gallici et al., 2024). While these techniques potentially side-step the issue, the consequences of a semi-gradient method may still exist. In Appendix C.3.1, we present an empirical analysis to gauge when krope’s learning algorithm may be unstable. We find that while individual off-policy transitions can determine the instability of fixed-point and semi-gradient algorithms such as fqe, pairs of off-policy transitions can determine krope’s instability. Since we are unlikely to have control over the distribution over pairs of transitions in practice, we need to resort to fundamental changes to the algorithm. One potential modification to overcome this limitation is based on that by Feng et al. (2019). Their key insight is to leverage the Legendre-Fenchel transformation and replace the fixed-point loss function of semi-gradient methods with an equivalent expression that avoids semi-gradient learning (Rockafellar & Wets, 1998). However, a drawback with this approach is that the new learning objective is a minimax procedure, which can be challenging to optimize in practice. In future work, we will explore the viability of this approach to design a provably convergent version of krope.
@@ 6 Conclusion
Impact Statement
Our work is largely focused on studying fundamental RL research questions, and thus we do not see any immediate negative societal impacts.
@@ A.1 Bisimulation Metrics
In this section, we present background information on bisimulations and its associated metrics. Our proposed representation learning algorithm is a bisimulation-based algorithm. Bisimulation abstractions are those under which two states with identical reward functions and that lead to identical groups of next states under any action are classified as similar (Ferns et al., 2004, 2011; Ferns & Precup, 2014). Bisimulations are the strictest forms of abstractions. In practice, the exact bisimulation criterion is difficult to satisfy computationally and statistically. A more relaxed version of this notion is the $\pi$-bisimulation metrics. These metrics capture the similarity between two states such that two states are considered similar if they have identical expected reward functions and expected transitions to identical groups of next states under a fixed policy $\pi$ (Castro, 2019).
@@
Using the above metrics, prior works have introduced several representation learning algorithms to learn representations such that the distance between representations in latent space model the above distance metrics (Castro et al., 2022, 2023; Zhang et al., 2021; Kemertas & Aumentado-Armstrong, 2021; Pavse & Hanna, 2023a).
@@ A.2 Reproducing Kernel Hilbert Spaces
The core usage of the rkhs is to precisely characterize the nature of the krope kernel. In practice, we deal with neural network representations, which are embedded in Euclidean space. Therefore, our goal is to learn representations in Euclidean space that approximate the properties of representations in the rkhs. For more details on the rkhs, we refer readers to (Castro et al., 2023) and (Gretton et al., 2012).
@@ Appendix B Theoretical Results
In this section, we present the proofs of our main and supporting theoretical results. The first set of proofs in Section B.1 show that krope is a valid operator. While new to our work, the proofs follow those by (Castro et al., 2023). The next set of proofs in Section B.2 prove the stability of krope representations and are novel to our work. For presentation purposes, it will often be convenient to refer to a state-action pair as $x\in\mathcal{X}$ instead of $(s,a)$.
@@
Given the definition of the krope kernel, we now present an operator that converges to $k^{\pi_{e}}$:
Definition 7 (krope operator).
[…]
We now present the proofs demonstrating the validity of the krope operator. All the proofs in this sub-section model those by (Castro et al., 2023). The primary difference is that our operator is for state-actions instead of states.
@@
Before presenting Lemma 3, we define the distance metric $d_{\text{{krope}}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ induced by the krope kernel $k^{{\pi_{e}}}$ as follows:
@@ B.2.1 Supporting Theoretical Results
We first define ${\pi_{e}}$-bisimulation (Castro, 2019). Note that (Castro, 2019) considered only state abstractions, while we consider state-action abstractions. $\phi$ is considered a ${\pi_{e}}$-bisimulation abstraction if it induces a mapping between $\mathcal{X}$ and $\mathcal{X}^{\phi}$ such that for any $x,y\in\mathcal{X}$ such that $x,y\in\phi(x)$, we have:
@@
Once we have $\phi$, we consider the following to show Bellman completeness. Our proof closely follows the proof of Proposition 20 from (Chen & Jiang, 2019). First recall the definition of Bellman completeness from Definition 8: $\forall f\in\mathcal{F},\forall\mathcal{T}^{{\pi_{e}}}f\in\mathcal{G}$, $\sup_{f\in\mathcal{F}}\inf_{g\in\mathcal{G}}\|g-\mathcal{T}^{{\pi_{e}}}f\|_{\infty}=0$. Given that the smallest value $\forall f\in\mathcal{F},\forall\mathcal{T}^{{\pi_{e}}}f\in\mathcal{G}$, $\sup_{f\in\mathcal{F}}\inf_{g\in\mathcal{G}}\|g-\mathcal{T}^{{\pi_{e}}}f\|_{\infty}$ can take on is zero, we will prove our claim by showing that $\forall f\in\mathcal{F},\forall\mathcal{T}^{{\pi_{e}}}f\in\mathcal{G}$, $\sup_{f\in\mathcal{F}}\inf_{g\in\mathcal{G}}\|g-\mathcal{T}^{{\pi_{e}}}f\|_{\infty}$ is upper-bounded by zero when $\phi$ is a ${\pi_{e}}$-bisimulation.
@@
Proposition 2.21 from (Paulsen & Raghupathi, 2016) states that any kernel $k$ is positive semidefinite if it takes the form: $k(a,b)=\min\{a,b\}$ where $a,b\in[0,\infty)$.
@@
From (Gretton et al., 2012), we have the definition of MMD between two probability distributions $p,q$ given kernel $k$:
@@
By Proposition 2.21 in (Paulsen & Raghupathi, 2016), the matrix $K$ is positive semidefinite, so we only need to show that $K$ is full rank. WLOG assume that $0<x_{1}<x_{2}<\cdots<x_{n}$. We prove by induction on $n$. The base case with $n=1$ clearly holds. Suppose the claim holds for $n-1$ numbers. Now consider $n$ numbers. Let $\alpha=(\alpha_{1},\ldots,\alpha_{n})^{\top}\in^{n}$. It suffices to show that $K\alpha=0$ implies $\alpha=0$. We write $K$ in block matrix form as